OpenStudy (anonymous):

1. (7 + √3)(7-√3) 2. (6+3√2)(5-√2) 3. 4√125

OpenStudy (anonymous):

ok , you need to know this first: \[\LARGE (a\pm b)(c\pm d)=ac\pm ad \pm bc \pm bd\] I hope that' not too confusing.. In your case you have: \[\Large (7+\sqrt3)(7-\sqrt3)=7\cdot 7-7\sqrt3 +\sqrt3\cdot 7-\sqrt3\cdot \sqrt3 \] now what can you simplify there? :) (question nr. 1)

OpenStudy (anonymous):

\[\Large \Large (7+\sqrt3)(7-\sqrt3)=7\cdot 7\cancel{-7\sqrt3} \cancel{+\sqrt3\cdot 7}-\sqrt3\cdot \sqrt3 \]

OpenStudy (anonymous):

what has left? :)

OpenStudy (anonymous):

@joemost still there ? :)

OpenStudy (anonymous):

the first one is: \[\LARGE (7+\sqrt3)(7-\sqrt3)=7^2-(\sqrt3\;)^2=49-3=46\] the second one.. \[\Large (6+3\sqrt2)(5-\sqrt2)=30-6\sqrt2+15\sqrt2-3\cdot 2\] \[\Large =24+9\sqrt2\] third one... \[\LARGE 4\sqrt{125}=4\sqrt{5^2\cdot 5 }=4\sqrt{5^2}\cdot \sqrt5 \] finish it...

OpenStudy (anonymous):

ahh thanks, i got it :D

OpenStudy (anonymous):

My Pleasure xD